\begin{thebibliography}{}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Baroni et~al.(2014)Baroni, Dinu, and Kruszewski}]{baroni:2014}
Marco Baroni, Georgiana Dinu, and Germ\'{a}n Kruszewski. 2014.
\newblock Don't count, predict! a systematic comparison of context-counting vs.
  context-predicting semantic vectors.
\newblock In {\em Proceedings of the 52nd Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)\/}. pages 238--247.

\bibitem[{Blei et~al.(2003)Blei, Ng, and Jordan}]{Blei2003}
David~M. Blei, Andrew~Y. Ng, and Michael~I. Jordan. 2003.
\newblock {Latent Dirichlet Allocation}.
\newblock {\em Journal of Machine Learning Research\/} 3:993--1022.

\bibitem[{Bruni et~al.(2014)Bruni, Tran, and Baroni}]{Bruni:2014}
Elia Bruni, Nam~Khanh Tran, and Marco Baroni. 2014.
\newblock Multimodal distributional semantics.
\newblock {\em Journal of Artificial Intelligence Research\/} 49:1--47.

\bibitem[{Chen et~al.(2015)Chen, Xu, He, and Wang}]{chen:2015}
Tao Chen, Ruifeng Xu, Yulan He, and Xuan Wang. 2015.
\newblock Improving distributed representation of word sense via wordnet gloss
  composition and context clustering.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 2: Short Papers)\/}. pages 15--20.

\bibitem[{Chen et~al.(2014)Chen, Liu, and Sun}]{chen-liu-sun:2014}
Xinxiong Chen, Zhiyuan Liu, and Maosong Sun. 2014.
\newblock A unified model for word sense representation and disambiguation.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)\/}. pages 1025--1035.

\bibitem[{Cheng and Kartsaklis(2015)}]{chengJianpeng:2015}
Jianpeng Cheng and Dimitri Kartsaklis. 2015.
\newblock Syntax-aware multi-sense word embeddings for deep compositional
  models of meaning.
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing\/}. pages 1531--1542.

\bibitem[{Cheng et~al.(2015)Cheng, Wang, Wen, Yan, and Chen}]{Cheng:2015}
Jianpeng Cheng, Zhongyuan Wang, Ji-Rong Wen, Jun Yan, and Zheng Chen. 2015.
\newblock Contextual text understanding in distributional semantic space.
\newblock In {\em Proceedings of the 24th ACM International on Conference on
  Information and Knowledge Management\/}. pages 133--142.

\bibitem[{Finkelstein et~al.(2002)Finkelstein, Gabrilovich, Matias, Rivlin,
  Solan, Wolfman, and Ruppin}]{Finkelstein:2001}
Lev Finkelstein, Evgeniy Gabrilovich, Yossi Matias, Ehud Rivlin, Zach Solan,
  Gadi Wolfman, and Eytan Ruppin. 2002.
\newblock Placing search in context: The concept revisited.
\newblock {\em ACM Transactions on Information Systems\/} 20:116--131.

\bibitem[{Flekova and Gurevych(2016)}]{flekova-gurevych:2016:P16-1}
Lucie Flekova and Iryna Gurevych. 2016.
\newblock Supersense embeddings: A unified model for supersense interpretation,
  prediction, and utilization.
\newblock In {\em Proceedings of the 54th Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)\/}. pages 2029--2041.

\bibitem[{Ghannay et~al.(2016)Ghannay, Favre, Estève, and
  Camelin}]{Ghannay:2016}
Sahar Ghannay, Benoit Favre, Yannick Estève, and Nathalie Camelin. 2016.
\newblock Word embedding evaluation and combination.
\newblock In {\em Proceedings of the Tenth International Conference on Language
  Resources and Evaluation (LREC 2016)\/}.

\bibitem[{Hill et~al.(2015)Hill, Reichart, and Korhonen}]{Hill:2015}
Felix Hill, Roi Reichart, and Anna Korhonen. 2015.
\newblock Simlex-999: Evaluating semantic models with genuine similarity
  estimation.
\newblock {\em Computational Linguistics\/} 41:665--695.

\bibitem[{Hoffman et~al.(2010)Hoffman, Bach, and Blei}]{Hoffman:2010}
Matthew Hoffman, Francis~R. Bach, and David~M. Blei. 2010.
\newblock Online learning for latent dirichlet allocation.
\newblock In {\em Advances in Neural Information Processing Systems 23\/}.
  pages 856--864.

\bibitem[{Huang et~al.(2012)Huang, Socher, Manning, and Ng}]{Huang:2012}
Eric~H. Huang, Richard Socher, Christopher~D. Manning, and Andrew~Y. Ng. 2012.
\newblock Improving word representations via global context and multiple word
  prototypes.
\newblock In {\em Proceedings of the 50th Annual Meeting of the Association for
  Computational Linguistics: Long Papers - Volume 1\/}. pages 873--882.

\bibitem[{Iacobacci et~al.(2015)Iacobacci, Pilehvar, and
  Navigli}]{iacobacci:2015}
Ignacio Iacobacci, Mohammad~Taher Pilehvar, and Roberto Navigli. 2015.
\newblock Sensembed: Learning sense embeddings for word and relational
  similarity.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)\/}. pages 95--105.

\bibitem[{Jauhar et~al.(2015)Jauhar, Dyer, and Hovy}]{jauhar:2015}
Sujay~Kumar Jauhar, Chris Dyer, and Eduard Hovy. 2015.
\newblock Ontologically grounded multi-sense representation learning for
  semantic vector space models.
\newblock In {\em Proceedings of the 2015 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies\/}. pages 683--693.

\bibitem[{Levy and Goldberg(2014)}]{Levy:2014}
Omer Levy and Yoav Goldberg. 2014.
\newblock Neural word embedding as implicit matrix factorization.
\newblock In {\em Advances in Neural Information Processing Systems 27\/}.
  pages 2177--2185.

\bibitem[{Levy et~al.(2015)Levy, Goldberg, and Dagan}]{Levy:2015}
Omer Levy, Yoav Goldberg, and Ido Dagan. 2015.
\newblock Improving distributional similarity with lessons learned from word
  embeddings.
\newblock {\em Transactions of the Association for Computational Linguistics\/}
  3:211--225.

\bibitem[{Li and Jurafsky(2015)}]{li-jurafsky:2015}
Jiwei Li and Dan Jurafsky. 2015.
\newblock Do multi-sense embeddings improve natural language understanding?
\newblock In {\em Proceedings of the 2015 Conference on Empirical Methods in
  Natural Language Processing\/}. pages 1722--1732.

\bibitem[{Liu et~al.(2015{\natexlab{a}})Liu, Qiu, and Huang}]{LiuPengfei:2015}
Pengfei Liu, Xipeng Qiu, and Xuanjing Huang. 2015{\natexlab{a}}.
\newblock Learning context-sensitive word embeddings with neural tensor
  skip-gram model.
\newblock In {\em Proceedings of the 24th International Conference on
  Artificial Intelligence\/}. pages 1284--1290.

\bibitem[{Liu et~al.(2015{\natexlab{b}})Liu, Liu, Chua, and Sun}]{Liu:2015}
Yang Liu, Zhiyuan Liu, Tat-Seng Chua, and Maosong Sun. 2015{\natexlab{b}}.
\newblock Topical word embeddings.
\newblock In {\em AAAI Conference on Artificial Intelligence\/}. pages
  2418--2424.

\bibitem[{Luong et~al.(2013)Luong, Socher, and Manning}]{luong:2013}
Thang Luong, Richard Socher, and Christopher Manning. 2013.
\newblock Better word representations with recursive neural networks for
  morphology.
\newblock In {\em Proceedings of the Seventeenth Conference on Computational
  Natural Language Learning\/}. pages 104--113.

\bibitem[{Ma and Hovy(2016)}]{ma-hovy:2016:P16-1}
Xuezhe Ma and Eduard Hovy. 2016.
\newblock End-to-end sequence labeling via bi-directional lstm-cnns-crf.
\newblock In {\em Proceedings of the 54th Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)\/}. pages 1064--1074.

\bibitem[{Mikolov et~al.(2013{\natexlab{a}})Mikolov, Chen, Corrado, and
  Dean}]{MikolovCoRR2013}
Tomas Mikolov, Kai Chen, Greg Corrado, and Jeffrey Dean. 2013{\natexlab{a}}.
\newblock Efficient estimation of word representations in vector space.
\newblock {\em CoRR\/} abs/1301.3781.

\bibitem[{Mikolov et~al.(2013{\natexlab{b}})Mikolov, Sutskever, Chen, Corrado,
  and Dean}]{MikolovSCCD13nips}
Tomas Mikolov, Ilya Sutskever, Kai Chen, Gregory~S. Corrado, and Jeffrey Dean.
  2013{\natexlab{b}}.
\newblock Distributed representations of words and phrases and their
  compositionality.
\newblock In {\em Advances in Neural Information Processing Systems 26: 27th
  Annual Conference on Neural Information Processing Systems 2013\/}. pages
  3111--3119.

\bibitem[{Modi(2016)}]{modi:2016}
Ashutosh Modi. 2016.
\newblock Event embeddings for semantic script modeling.
\newblock In {\em Proceedings of the Conference on Computational Natural
  Language Learning\/}. pages 75--83.

\bibitem[{Modi and Titov(2014)}]{modititov:2014}
Ashutosh Modi and Ivan Titov. 2014.
\newblock Inducing neural models of script knowledge.
\newblock In {\em Proceedings of the Eighteenth Conference on Computational
  Natural Language Learning\/}. pages 49--57.

\bibitem[{Modi et~al.(2017)Modi, Titov, Demberg, Sayeed, and
  Pinkal}]{ashutoshTacl2017}
Ashutosh Modi, Ivan Titov, Vera Demberg, Asad Sayeed, and Manfred Pinkal. 2017.
\newblock Modelling semantic expectation: Using script knowledge for referent
  prediction.
\newblock {\em Transactions of the Association for Computational Linguistics\/}
  5:31--44.

\bibitem[{Neelakantan et~al.(2014)Neelakantan, Shankar, Passos, and
  McCallum}]{neelakantan:2014}
Arvind Neelakantan, Jeevan Shankar, Alexandre Passos, and Andrew McCallum.
  2014.
\newblock Efficient non-parametric estimation of multiple embeddings per word
  in vector space.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP)\/}. pages 1059--1069.

\bibitem[{Nguyen et~al.(2015{\natexlab{a}})Nguyen, Billingsley, Du, and
  Johnson}]{nguyen2015}
Dat~Quoc Nguyen, Richard Billingsley, Lan Du, and Mark Johnson.
  2015{\natexlab{a}}.
\newblock {Improving Topic Models with Latent Feature Word Representations}.
\newblock {\em Transactions of the Association for Computational Linguistics\/}
  3:299--313.

\bibitem[{Nguyen et~al.(2017)Nguyen, Dras, and Johnson}]{jPTDP2017}
Dat~Quoc Nguyen, Mark Dras, and Mark Johnson. 2017.
\newblock {A Novel Neural Network Model for Joint POS Tagging and Graph-based
  Dependency Parsing}.
\newblock In {\em Proceedings of the CoNLL 2017 Shared Task: Multilingual
  Parsing from Raw Text to Universal Dependencies\/}.

\bibitem[{Nguyen et~al.(2015{\natexlab{b}})Nguyen, Sirts, and
  Johnson}]{NguyenALTA2015}
Dat~Quoc Nguyen, Kairit Sirts, and Mark Johnson. 2015{\natexlab{b}}.
\newblock {Improving Topic Coherence with Latent Feature Word Representations
  in MAP Estimation for Topic Modeling}.
\newblock In {\em Proceedings of the Australasian Language Technology
  Association Workshop 2015\/}. pages 116--121.

\bibitem[{Pennington et~al.(2014)Pennington, Socher, and
  Manning}]{Pennington2014}
Jeffrey Pennington, Richard Socher, and Christopher~D. Manning. 2014.
\newblock Glove: Global vectors for word representation.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing (EMNLP 2014)\/}. pages 1532--1543.

\bibitem[{Pilehvar and Collier(2016)}]{PilehvarC16}
Mohammad~Taher Pilehvar and Nigel Collier. 2016.
\newblock De-conflated semantic representations.
\newblock In {\em Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing\/}. pages 1680--1690.

\bibitem[{Qiu et~al.(2014)Qiu, Cui, Bian, Gao, and Liu}]{qiu-EtAl:2014}
Siyu Qiu, Qing Cui, Jiang Bian, Bin Gao, and Tie-Yan Liu. 2014.
\newblock Co-learning of word representations and morpheme representations.
\newblock In {\em Proceedings of COLING 2014, the 25th International Conference
  on Computational Linguistics: Technical Papers\/}. pages 141--150.

\bibitem[{Rastogi et~al.(2015)Rastogi, Van~Durme, and Arora}]{rastogi2015NAACL}
Pushpendre Rastogi, Benjamin Van~Durme, and Raman Arora. 2015.
\newblock {Multiview LSA: Representation Learning via Generalized CCA}.
\newblock In {\em Proceedings of the 2015 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies\/}. pages 556--566.

\bibitem[{{\v R}eh{\r u}{\v r}ek and Sojka(2010)}]{rehurek_lrec}
Radim {\v R}eh{\r u}{\v r}ek and Petr Sojka. 2010.
\newblock {Software Framework for Topic Modelling with Large Corpora}.
\newblock In {\em {Proceedings of the LREC 2010 Workshop on New Challenges for
  NLP Frameworks}\/}. pages 45--50.

\bibitem[{Reisinger and Mooney(2010)}]{Reisinger:2010}
Joseph Reisinger and Raymond~J. Mooney. 2010.
\newblock Multi-prototype vector-space models of word meaning.
\newblock In {\em Human Language Technologies: The 2010 Annual Conference of
  the North American Chapter of the Association for Computational
  Linguistics\/}. pages 109--117.

\bibitem[{Rothe and Sch{\"{u}}tze(2015)}]{RotheS15}
Sascha Rothe and Hinrich Sch{\"{u}}tze. 2015.
\newblock Autoextend: Extending word embeddings to embeddings for synsets and
  lexemes.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics and the 7th International Joint Conference on
  Natural Language Processing, Volume 1: Long Papers\/}. pages 1793--1803.

\bibitem[{Schnabel et~al.(2015)Schnabel, Labutov, Mimno, and
  Joachims}]{schnabel2015eval}
Tobias Schnabel, Igor Labutov, David Mimno, and Thorsten Joachims. 2015.
\newblock Evaluation methods for unsupervised word embeddings.
\newblock In {\em Proceedings of the Conference on Empirical Methods in Natural
  Language Processing\/}. pages 298--307.

\bibitem[{Schwartz et~al.(2015)Schwartz, Reichart, and
  Rappoport}]{schwartz:2015}
Roy Schwartz, Roi Reichart, and Ari Rappoport. 2015.
\newblock Symmetric pattern based word embeddings for improved word similarity
  prediction.
\newblock In {\em Proceedings of CoNLL 2015\/}. pages 258--267.

\bibitem[{Shaoul and Westbury(2010)}]{shaoul2010westbury}
Cyrus Shaoul and Chris Westbury. 2010.
\newblock The westbury lab wikipedia corpus.
\newblock {\em Edmonton, AB: University of Alberta\/} .

\bibitem[{Socher et~al.(2013)Socher, Perelygin, Wu, Chuang, Manning, Ng, and
  Potts}]{socher-EtAl:2013}
Richard Socher, Alex Perelygin, Jean Wu, Jason Chuang, Christopher~D. Manning,
  Andrew Ng, and Christopher Potts. 2013.
\newblock Recursive deep models for semantic compositionality over a sentiment
  treebank.
\newblock In {\em Proceedings of the 2013 Conference on Empirical Methods in
  Natural Language Processing\/}. pages 1631--1642.

\bibitem[{Sutskever et~al.(2014)Sutskever, Vinyals, and Le}]{Sutskever:2014}
Ilya Sutskever, Oriol Vinyals, and Quoc~V. Le. 2014.
\newblock Sequence to sequence learning with neural networks.
\newblock In {\em Proceedings of the 27th International Conference on Neural
  Information Processing Systems\/}. pages 3104--3112.

\bibitem[{Tian et~al.(2014)Tian, Dai, Bian, Gao, Zhang, Chen, and
  Liu}]{tian-EtAl:2014}
Fei Tian, Hanjun Dai, Jiang Bian, Bin Gao, Rui Zhang, Enhong Chen, and Tie-Yan
  Liu. 2014.
\newblock A probabilistic model for learning multi-prototype word embeddings.
\newblock In {\em Proceedings of COLING 2014, the 25th International Conference
  on Computational Linguistics: Technical Papers\/}. pages 151--160.

\bibitem[{Vilnis and McCallum(2015)}]{vilnis2015}
Luke Vilnis and Andrew McCallum. 2015.
\newblock Word representations via gaussian embedding.
\newblock {\em International Conference on Learning Representations (ICLR)\/} .

\bibitem[{Wu and Giles(2015)}]{Wu:2015}
Zhaohui Wu and C.~Lee Giles. 2015.
\newblock Sense-aware semantic analysis: A multi-prototype word representation
  model using wikipedia.
\newblock In {\em Proceedings of the Twenty-Ninth AAAI Conference on Artificial
  Intelligence\/}. pages 2188--2194.

\bibitem[{Zhang and Zhong(2016)}]{Zhang2016}
Heng Zhang and Guoqiang Zhong. 2016.
\newblock Improving short text classification by learning vector
  representations of both words and hidden topics.
\newblock {\em Knowledge-Based Systems\/} 102:76--–86.

\end{thebibliography}

\documentclass[11pt,a4paper]{article}
\usepackage[nohyperref]{acl2017}
\usepackage{times}
\usepackage{latexsym}

\usepackage{url}

\usepackage{amsmath}
\usepackage{multirow}
\usepackage{url}
\usepackage{graphicx}
\usepackage[ruled,vlined]{algorithm2e}
\usepackage{amssymb}
\usepackage{amsfonts}
\usepackage{todonotes}
\usepackage{fancyvrb}
\usepackage{tikz}
\aclfinalcopy 

\title{A Mixture Model for Learning Multi-Sense Word Embeddings}

\author{
Dai Quoc Nguyen${}^{1}$, Dat Quoc Nguyen${}^{2}$, Ashutosh Modi${}^{1}$, Stefan Thater${}^{1}$, Manfred Pinkal${}^{1}$\\
\\
${}^{1}$Department of Computational Linguistics, Saarland University, Germany\\
{\tt{{\{daiquocn, ashutosh, stth, pinkal\}@coli.uni-saarland.de}}} \\
${}^{2 }$Department of Computing, Macquarie University, Australia \\
{\tt{{dat.nguyen@students.mq.edu.au}}}
}

\begin{document}
\maketitle

\begin{abstract}

Word embeddings are now a standard technique for inducing meaning representations for words. For getting good representations, it is important to take into account different senses of a word. In this paper, we propose a mixture model for learning multi-sense word embeddings. Our model generalizes the previous works in that it allows to induce different weights of different senses of a word. The experimental results show that our model outperforms previous models on standard evaluation tasks.
 
\end{abstract}

\section{Introduction}
Word embeddings have shown to be useful in various NLP tasks such as sentiment analysis, topic models, script learning, machine translation, sequence labeling and parsing \cite{socher-EtAl:2013,Sutskever:2014,modititov:2014,nguyen2015,NguyenALTA2015,modi:2016,ma-hovy:2016:P16-1,jPTDP2017,ashutoshTacl2017}.
A word embedding captures the syntactic and semantic properties of a word by representing the word in a form of a real-valued vector \cite{MikolovCoRR2013,MikolovSCCD13nips,Pennington2014,Levy:2014}. 

However, usually word embedding models do not take into account lexical ambiguity. 
For example, the word $bank$ is usually represented by a single vector representation for all senses including \textit{sloping land} and \textit{financial institution}.
Recently, approaches have been proposed to learn multi-sense word embeddings, where each sense of a word corresponds to a sense-specific embedding.
\newcite{Reisinger:2010}, \newcite{Huang:2012} and \newcite{Wu:2015} proposed methods to cluster the contexts of each word and then using cluster centroids as vector representations for word senses. 
\newcite{neelakantan:2014}, \newcite{tian-EtAl:2014}, \newcite{li-jurafsky:2015} and \newcite{chen:2015} extended  Word2Vec models \cite{MikolovCoRR2013,MikolovSCCD13nips} to learn a vector representation for each sense of a word. 
\newcite{chen-liu-sun:2014}, \newcite{iacobacci:2015} and \newcite{flekova-gurevych:2016:P16-1} performed word sense induction using external resources (e.g., WordNet, BabelNet) and then learned sense embeddings using the Word2Vec models. 
\newcite{RotheS15} and \newcite{PilehvarC16} presented methods using pre-trained word embeddings to learn embeddings from WordNet synsets.
\newcite{Cheng:2015}, \newcite{Liu:2015}, \newcite{LiuPengfei:2015} and \newcite{Zhang2016} directly opt the Word2Vec Skip-gram model \cite{MikolovSCCD13nips} for learning the embeddings of words and topics on a topic-assigned corpus.

One issue in these previous works is that they assign the same weight to every sense of a word.
The central assumption of our work is that each sense of a word given a context, should correspond to a mixture of weights reflecting  different association degrees of the word  with multiple senses in the context.
The mixture weights will help to model word meaning better.

In this paper, we propose a new model for learning \textbf{M}ulti-\textbf{S}ense \textbf{W}ord \textbf{E}mbeddings (\textsc{mswe}). 
Our \textsc{mswe} model learns vector representations of a word based on a mixture of its sense representations.  
The key difference between \textsc{mswe} and other models is that we induce the weights of senses while jointly learning the word and sense embeddings.
Specifically, we train a topic model \citep{Blei2003} to obtain the topic-to-word and document-to-topic probability distributions which are then used to infer the weights of topics. We use these weights to define a compositional vector representation for each target word to predict its context words. \textsc{mswe} thus is different from the topic-based models \citep{Cheng:2015,Liu:2015,LiuPengfei:2015,Zhang2016}, in which we do not use the topic assignments when jointly learning vector representations of words and topics. Here we not only learn vectors based on the most suitable topic of a word given its context, but  we also take into consideration all possible meanings of the word.   

The main contributions of our study are:
(i) We introduce a mixture model  for learning word and sense embeddings (\textsc{mswe}) by inducing mixture weights of word senses.
(ii) We show that \textsc{mswe} performs better than the baseline Word2Vec Skip-gram 
and other embedding models on the word analogy task \cite{MikolovCoRR2013} and the word similarity task \cite{Reisinger:2010}.

\section{The mixture model}
\label{sec:mswe}

In this section, we present the mixture model for learning multi-sense word embeddings. 
Here we treat topics as senses. The model learns a representation for each word using a mixture of its topical representations.

Given a number of topics and a corpus $D$ of documents $d = \{w_{d,1}, w_{d,2},..., w_{d,M_d}\}$, we apply a topic model \cite{Blei2003} to obtain the topic-to-word $\Pr(w|t)$ and document-to-topic $\Pr(t|d)$ probability distributions. We then infer a weight for the $m^{th}$ word $w_{d,m}$ with topic $t$ in document $d$:
\begin{equation}
\lambda_{d,m,t} = \Pr(w_{d,m}|t) \times \Pr(t|d)
\label{equal:weight}
\end{equation}

We define two \textsc{mswe} variants: \textsc{mswe-1} learns vectors for words based on the most suitable topic given  document $d$  while \textsc{mswe-2} marginalizes over all senses of a word to take into account all possible senses of the word:
\begin{eqnarray}
\textsc{mswe-1:} & \boldsymbol{s}_{w_{d,m}} &= \dfrac{\boldsymbol{v}_{w_{d,m}} + \lambda_{d,m,t'} \times \boldsymbol{v}_{t'}}{1 + \lambda_{d,m,t'}} \nonumber \\
\textsc{mswe-2:} & \boldsymbol{s}_{w_{d,m}} &= \dfrac{\boldsymbol{v}_{w_{d,m}} + \sum_{t=1}^{T} \lambda_{d,m,t} \times \boldsymbol{v}_{t}}{1 + \sum_{t=1}^{T} \lambda_{d,m,t}} \nonumber
\end{eqnarray}
where $\boldsymbol{s}_{w_{d,m}}$ is the compositional vector representation of the $m^{th}$ word $w_{d,m}$ and the topics in document $d$; $\boldsymbol{v}_w$ is the target vector representation of a word type $w$ in vocabulary $V$; $\boldsymbol{v}_{t}$ is the vector representation of topic $t$; $T$ is the number of topics; $\lambda_{d,m,t}$ is defined as in Equation \ref{equal:weight}, and in \textsc{mswe-1} we define $t' = \underset{t}{\arg\max} \ \lambda_{d,m,t}$.  

We learn representations by minimizing the following negative log-likelihood function:
\begin{equation}
\resizebox{.89\hsize}{!}{$
\mathcal{L}  = - \sum\limits_{d \in D} \sum\limits_{m=1}^{M_d} \sum\limits_{\substack{-k \leq j \leq k \\ j \neq 0}} \log \Pr(\tilde{\boldsymbol{v}}_{w_{d,m+j}}|\boldsymbol{s}_{w_{d,m}})
$}
\label{equal:1}
\end{equation}

\noindent where the $m^{th}$ word $w_{d,m}$ in document $d$ is a target word while the $(m+j)^{th}$ word $w_{d,m+j}$ in document  $d$  is a context word of $w_{d,m}$ and $k$ is the context size. In addition, $\tilde{\boldsymbol{v}}_{w}$ is the context vector representation of the word type $w$. 
The probability $\Pr(\tilde{\boldsymbol{v}}_{w_{d,m+j}}|\boldsymbol{s}_{w_{d,m}})$ is defined using the softmax function as follows:
\begin{equation*}
\Pr(\tilde{\boldsymbol{v}}_{w_{d,m+j}}|\boldsymbol{s}_{w_{d,m}}) = \frac{\exp(\tilde{\boldsymbol{v}}_{w_{d,m+j}}^\mathsf{T} \boldsymbol{s}_{w_{d,m}})}{\sum_{c' \in V} \exp(\tilde{\boldsymbol{v}}_{c'}^\mathsf{T} \boldsymbol{s}_{w_{d,m}})} 
\label{equal:2}
\end{equation*}

Since computing $\log\Pr(\tilde{\boldsymbol{v}}_{w_{d,m+j}}|\boldsymbol{s}_{w_{d,m}})$ is expensive for each training instance, we approximate $\log\Pr(\tilde{\boldsymbol{v}}_{w_{d,m+j}}|\boldsymbol{s}_{w_{d,m}})$ in Equation \ref{equal:1} with the following negative-sampling objective \cite{MikolovSCCD13nips}:
\begin{eqnarray}
\mathcal{O}_{d,m,m+j}& = &\log \sigma \left(\tilde{\boldsymbol{v}}_{w_{d,m+j}}^\mathsf{T} \boldsymbol{s}_{w_{d,m}}\right) \nonumber \\ &+& \sum_{i=1}^{K} \log \sigma \left(-\tilde{\boldsymbol{v}}_{c_i}^\mathsf{T} \boldsymbol{s}_{w_{d,m}}\right)
\label{equal:3}
\end{eqnarray} 

\noindent where 
each word $c_i$ is sampled from a noise distribution.\footnote{We use an unigram distribution raised to the 3/4 power \cite{MikolovSCCD13nips} as the noise distribution.}
In fact, \textsc{mswe} can be viewed as a generalization of the well-known  Word2Vec Skip-gram model with negative sampling \cite{MikolovSCCD13nips} where all the mixture weights $\lambda_{d,m,t}$ are set to zero. The models are trained using Stochastic Gradient Descent (SGD).


\section{Experiments}
\label{sec:exp}

We evaluate \textsc{mswe} on two different tasks: word similarity and word analogy. We also provide experimental results obtained by the baseline Word2Vec Skip-gram model and other previous works. 

Note that not all previous results are mentioned in this paper for comparison because the training corpora used in most previous research work are much larger than ours \citep{baroni:2014,li-jurafsky:2015,schwartz:2015,Levy:2015}. Also there are differences in the pre-processing steps that could affect the results. 
We could also improve obtained results by using a larger training corpus, but this is not central point of our paper. The objective of our paper is that the embeddings of topic and word can be combined into a single mixture model, leading to good improvements as established empirically.

\subsection{Experimental Setup}

Following \citet{Huang:2012} and \citet{neelakantan:2014}, we use the  Wesbury Lab Wikipedia corpus \citep{shaoul2010westbury} containing over 2M articles with about 990M  words for training.
In the preprocessing step, texts are lowercased and tokenized, numbers are mapped to 0, and punctuation marks are removed.
We extract a vocabulary of  200,000 most frequent word tokens  from the pre-processed corpus.
Words not occurring in the vocabulary are mapped to a special token \textsc{unk}, in which we use the embedding of \textsc{unk} for unknown words in the benchmark datasets.

We firstly use a small subset extracted from the \textsc{ws353} dataset \citep{Finkelstein:2001} to tune the hyper-parameters of the baseline Word2Vec Skip-gram model  for the word similarity task (see Section \ref{wordsim} for the task definition). We then directly use the tuned hyper-parameters for our \textsc{mswe} variants.
Vector size is also a hyper-parameter. While some approaches use a higher number of dimensions to obtain better results, we fix the vector size to be 300 as used by the baseline for a fair comparison.
The vanilla Latent Dirichlet Allocation (LDA) topic model \citep{Blei2003} is not scalable to a very large corpus, so we explore faster online topic models developed for large corpora. 
We train the online LDA topic  model \citep{Hoffman:2010} on the training corpus, and use the output of this topic model to compute the mixture weights as in Equation \ref{equal:weight}.\footnote{We use default parameters in \textit{gensim} \protect{\citep{rehurek_lrec}} for the online LDA model.} We also use the same \textsc{ws353} subset to tune the numbers of topics $T \in \{50, 100, 200, 300, 400\}$. We find that the most suitable numbers are $T = 50$ and $T = 200$ then used for all our experiments.
Here we learn 300-dimensional embeddings with the fixed context size $k=5$ (in Equation \ref{equal:1}) and $K = 10$ (in Equation \ref{equal:3}) as used by the baseline. 
During training, we randomly initialize model parameters (i.e. word and topic embeddings) and then learn them by using SGD with the initial learning rate of 0.01.

\subsection{Word Similarity}\label{wordsim}

The word similarity task evaluates the quality of word embedding models \cite{Reisinger:2010}. For a given dataset of word pairs, the evaluation is done by calculating correlation between the similarity scores of corresponding word embedding pairs with the human judgment scores. Higher Spearman's rank correlation ($\rho$) reflects better word embedding model. 
We evaluate \textsc{mswe} on standard datasets (as given in Table \ref{tab:datasets}) for the word similarity evaluation task.

\setlength{\abovecaptionskip}{3pt plus 1pt minus 1pt} 
\begin{table}[!t]
\centering
\setlength{\tabcolsep}{0.25em}
\resizebox{7.5cm}{!}{
\begin{tabular}{l|c|l}
\hline
Dataset & Word pairs & Reference \\
\hline
\textsc{ws353} & 353 & \newcite{Finkelstein:2001} \\
\textsc{SimLex} & 999 &  \newcite{Hill:2015} \\
\textsc{scws} & 2003 & \newcite{Huang:2012} \\
\textsc{rw} & 2034 & \newcite{luong:2013}\\
\textsc{men} & 3000 & \newcite{Bruni:2014} \\
\hline
\end{tabular}
}
\caption{The benchmark datasets. \textsc{ws353}: WordSimilarity-353. \textsc{rw}: Rare-Words. \textsc{SimLex}: SimLex-999. \textsc{scws}:  Stanford's Contextual Word Similarities. \textsc{men}: The MEN Test Collection. Each dataset contains similarity scores of human judgments for pairs of words.}
\label{tab:datasets}
\end{table}

Following \citet{Reisinger:2010}, \citet{Huang:2012}, \citet{neelakantan:2014}, we compute the similarity scores for a pair of words $\left(w, w'\right)$ with or without their respective contexts $\left(c, c'\right)$ as:

\vspace{-5pt}

{\small
\begin{eqnarray}
&&GlobalSim\left(w, w'\right) = \cos\left(\boldsymbol{v}_{w}, \boldsymbol{v}_{w'}\right)  \nonumber\\
&&AvgSim\left(w, w'\right) = \frac{1}{T^2} \sum_{t=1}^{T} \sum_{t'=1}^{T} \cos\left(\boldsymbol{v}_{w,t}, \boldsymbol{v}_{w',t'}\right) \nonumber\\
&& AvgSimC\left(w, w'\right) \nonumber \\ &&= \frac{1}{T^2} \sum_{t=1}^{T} \sum_{t'=1}^{T} \Big(\delta\left(\boldsymbol{v}_{w,t}, \boldsymbol{v}_{c}\right) \times \delta\left(\boldsymbol{v}_{w',t'}, \boldsymbol{v}_{c'}\right)  \nonumber \\
&&  \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \  \times \cos\left(\boldsymbol{v}_{w,t}, \boldsymbol{v}_{w',t'}\right)\Big) \nonumber
\end{eqnarray}
}

\vspace{-5pt}

\noindent where $\boldsymbol{v}_{w}$ is the vector representation of the word $w$, 
$\boldsymbol{v}_{w,t}$ is the multiple representation of the word $w$ and the topic $t$,
$\boldsymbol{v}_c$ is the vector representation of the context $c$ of the word $w$.
And $\cos\left(\boldsymbol{v}, \boldsymbol{v'}\right)$ is the cosine similarity between two vectors $\boldsymbol{v}$ and $\boldsymbol{v'}$. 
For our experiments, we set $\boldsymbol{v}_{w,t} = \boldsymbol{v}_{w} \oplus \left(\Pr(w|t) \times \boldsymbol{v}_{t}\right)$ and 
 $\boldsymbol{v}_c = \left(\frac{1}{|c|} \sum_{w \in c} \boldsymbol{v}_w\right) \oplus \left(\sum_{t} \Pr\left(t|c\right) \times \boldsymbol{v}_t\right)$, in which $\oplus$ is the concatenation operation and 
$\Pr\left(t|c\right)$ is inferred from the topic models by considering context $c$ as a document.
$GlobalSim$ only regards word embeddings, while $AvgSim$ considers multiple representations to capture different meanings (i.e. topics) and usages of a word.
$AvgSimC$ generalizes $AvgSim$ by taking into account the likelihood $\delta\left(\boldsymbol{v}_{w,t}, \boldsymbol{v}_{c}\right)$ that word $w$ takes topic $t$ given context $c$. $\delta\left(\boldsymbol{v}, \boldsymbol{v'}\right)$ is the inverse of the cosine distance from $\boldsymbol{v}$ to $\boldsymbol{v'}$ \cite{Huang:2012,neelakantan:2014}.  
{
\begin{table}[!t]
\centering
\setlength{\tabcolsep}{0.25em}
\resizebox{7.8cm}{!}{
\begin{tabular}{l|l|l|l|l|l}
\hline
Model & \textsc{rw} & \textsc{SimLex} & \textsc{scws} & \textsc{ws353} & \textsc{men} \\
\hline
\hline
\newcite{Huang:2012} & -- & -- & 58.6 & 71.3 & -- \\
\newcite{luong:2013}  & 34.36 & -- & 48.48  &  64.58 & --\\ 
\newcite{qiu-EtAl:2014} & 32.13 & -- & 53.40 & 65.19 & -- \\
\newcite{neelakantan:2014} & -- & -- & 65.5 &  69.2 & --\\
\newcite{chen-liu-sun:2014} & -- & -- & 64.2 & -- & --\\
\newcite{Hill:2015} & -- & \textbf{41.4} & -- & 65.5 & 69.9\\
\newcite{vilnis2015} & -- & 32.23 & -- & 65.49 & 71.31\\
\newcite{schnabel2015eval} & -- & -- & -- & 64.0 & 70.7 \\
\newcite{rastogi2015NAACL} & 32.9 & 36.7 & 65.6 & 70.8 & 73.9\\
\newcite{flekova-gurevych:2016:P16-1} & -- & -- & -- & -- & 74.26 \\
\hline
\hline
Word2Vec Skip-gram  & 32.64 & 38.20 &66.37 & {71.61} & 75.49 \\
\hline
$\textsc{mswe-1}_{50}$ & 34.85 & 38.77 & \textbf{66.83} & \textbf{72.40} & \underline{76.23} \\
$\textsc{mswe-1}_{200}$  & \underline{35.27} & 38.70 &  \underline{66.80} & 72.05 & 76.05 \\
\hline
$\textsc{mswe-2}_{50}$ & 34.98 & 38.79 & 66.61 & 71.71 & 75.90 \\
$\textsc{mswe-2}_{200}$  & \textbf{35.56}$^\star$ & \underline{39.19}$^\star$ & 66.65 & \underline{72.29} & \textbf{76.37}$^\star$ \\
\hline
\hline
\end{tabular}
}
\caption{Spearman's rank correlation ($\rho \times 100$) for the word similarity task when using $GlobalSim$. Subscripts 50 and 200 denote the online LDA topic model trained with $T=50$ and $T=200$ topics, respectively. 
$^\star$ denotes that our best score is significantly higher than the score of the baseline (with $p <$ 0.05, online toolkit from \url{http://www.philippsinger.info/?p=347}). Scores in \textbf{bold} and \underline{underline} are the best and second best scores.}
\label{tab:wordsimilarity}
\end{table}
}

\subsubsection{Results for word similarity}
\label{sec:wordsim}

Table \ref{tab:wordsimilarity} compares the evaluation results of \textsc{mswe} with results reported in prior work on the standard word similarity task when using $GlobalSim$.
We use subscripts 50 and 200 to denote the topic model trained with $T=50$ and $T=200$ topics, respectively.
Table \ref{tab:wordsimilarity} shows that our model  outperforms the baseline Word2Vec Skip-gram model (in fifth row from bottom).
Specifically, on the \textsc{rw} dataset, \textsc{mswe} obtains a significant improvement of $2.92$
in the Spearman's rank correlation (which is about 8.5\% relative improvement).

Compared to the published results, \textsc{mswe} obtains the highest accuracy on the \textsc{rw}, \textsc{scws}, \textsc{ws353} and \textsc{men} datasets, and achieves the second highest result on the \textsc{SimLex} dataset.
These indicate that \textsc{mswe} learns better representations for words taking into account different meanings.


{
\begin{table}[!t]
\centering
\setlength{\tabcolsep}{0.25em}
\resizebox{7.5cm}{!}{
\begin{tabular}{l|c|c}
\hline
Model & AvgSim & AvgSimC \\
\hline
\hline
\newcite{Huang:2012}  & 62.8  & 65.7   \\
\newcite{neelakantan:2014}   & \textbf{67.3}  & \textbf{69.3}  \\
\newcite{chen-liu-sun:2014}  & 66.2   & \underline{68.9}  \\
\newcite{chen:2015} & 65.7  & 66.4  \\
\newcite{Wu:2015} & -- & 66.4  \\
\newcite{jauhar:2015} & -- & 65.7  \\
\newcite{chengJianpeng:2015} & 62.5   & -- \\
\newcite{iacobacci:2015} & 62.4   & --\\
\newcite{Cheng:2015} & -- & 65.9 \\
\hline
\hline
$\textsc{mswe-1}_{50}$ & 66.6  & {66.7} \\
$\textsc{mswe-1}_{200}$ & \underline{66.7} & 66.6 \\
\hline
$\textsc{mswe-2}_{50}$ & 66.4  & 66.6 \\
$\textsc{mswe-2}_{200}$ & 66.6  & 66.6 \\
\hline
\hline
\end{tabular}
}
\caption{Spearman's rank correlation ($\rho \times 100$) on   \textsc{scws}, using $AvgSim$ and $AvgSimC$.}
\label{tab:scws}
\end{table}
}

\subsubsection{Results for contextual word similarity}
We evaluate our model \textsc{mswe} by using $AvgSim$ and $AvgSimC$ on the benchmark \textsc{scws} dataset which considers effects of the contextual information on the word similarity task.
As shown in Table \ref{tab:scws}, \textsc{mswe}  scores better than the closely related model proposed by \newcite{Cheng:2015} and generally obtains good results for this context sensitive dataset.
Although we produce better scores than \newcite{neelakantan:2014} and \newcite{chen-liu-sun:2014} when using $GlobalSim$, we are outperformed by them when using $AvgSim$ and $AvgSimC$. 
\newcite{neelakantan:2014} clustered the embeddings of the context words around each target word to predict its sense and \newcite{chen-liu-sun:2014} used pre-trained word embeddings to initialize vector representations of senses taken from WordNet, while we use a fixed number of topics as senses for words in \textsc{mswe}.


\subsection{Word Analogy}

{
\begin{table}[!t]
\centering
\begin{tabular}{l|c}
\hline
Model & Accuracy (\%)\\
\hline
\hline
\newcite{Pennington2014} & \textbf{70.3} \\
\newcite{baroni:2014} & 68.0 \\
\newcite{neelakantan:2014} & 64.0 \\
\newcite{Ghannay:2016} & 62.3 \\
\hline
\hline
Word2Vec Skip-gram  & 68.6 \\
\hline
\hline
$\textsc{mswe-1}_{50}$ & 69.6 \\
$\textsc{mswe-1}_{200}$ & \underline{69.9} \\
\hline
$\textsc{mswe-2}_{50}$ & 69.7 \\
$\textsc{mswe-2}_{200}$ & 69.5 \\
\hline
\hline
\end{tabular}
\caption{Accuracies  for the word analogy task. 
All our results are significantly higher than the result of Word2Vec Skip-gram (with two-tail $p < 0.001$ using  McNemar's test). 
\newcite{Pennington2014} used a larger training corpus of 1.6B words.
}
\label{tab:wordanalogy}
\end{table}
}

We evaluate the embedding models on the word analogy task introduced by \newcite{MikolovCoRR2013}. 
The task aims to answer questions in the form of ``$a$ is to $b$ as $c$ is to \_ $?$'', denoted as ``\emph{a : b $\rightarrow$ c : ?}'' (e.g., ``\emph{Hanoi : Vietnam  $\rightarrow$ Bern : ?}'').
There are 8,869 semantic and 10,675 syntactic questions grouped into 14 categories.
Each question is answered by finding the most suitable word closest to ``\emph{$\boldsymbol{v}_{b} - \boldsymbol{v}_{a} + \boldsymbol{v}_{c}$}'' measured by the cosine similarity.
The answer is correct only if the found closest word is exactly the same as the gold-standard (correct) one for the question. 

We report accuracies in Table \ref{tab:wordanalogy} and show that \textsc{mswe} achieves better results in comparison with the baseline Word2Vec Skip-gram. 
In particular, \textsc{mswe} reaches the accuracies of around 69.7$\%$ which is higher than the accuracy of 68.6$\%$ obtained by Word2Vec Skip-gram.

\section{Conclusions}
\label{sec:conclusion}

In this paper, we described a mixture model for learning multi-sense embeddings. 
Our model induces mixture weights to represent a word given context based on a mixture of its sense representations. 
The results show that our model scores better than Word2Vec, and produces highly competitive results on the standard evaluation tasks.  
In future work, we will explore better methods for taking into account the contextual information. 
We also plan to explore different approaches to compute the mixture weights in our model.
For example, if there is a large sense-annotated corpus available for training, the mixture weights could be defined based on the frequency (sense-count) distributions, instead of using the probability distributions produced by a topic model.
Furthermore, it is possible to consider the weights of senses as additional model parameters to be then learned during training.

\section*{Acknowledgments}
This research was funded by the German Research Foundation (DFG) as part of SFB 1102 ``Information Density and Linguistic Encoding''.
We would like to thank anonymous reviewers for their helpful comments.

\bibliographystyle{acl_natbib}
\bibliography{references}

\end{document}